(a) Find the Norton equivalent circuit of the circuit below after increasing the magnitude of the voltage source to 38 V. 52(a) Find the Norton equivalent circuit of the circuit below after increasing the magnitude of the voltage source to 38 V.
Answer #1

a) finding In (or) short circuit a and b current between more 104 va un 20-4 Ivo bi Ix a un 38V ♡ 40 qfx » $80 write kch at N= 38-va Vb-38 bave 3 - 2 ( 38-Na 20 Six = 38_Na Ubygg easy us ++ - - - - us (3*2*3*2) - V4 (+ + 3)-75730 V5 236 - 5 ho 35 v$.: IN= Vb - 201829 = 10.4148 .: INE 20.829 = 1044A IN = 10.414 A finding of RN short circuit the 381 source and connect a carwrite kaf at vx:- - o | vw (+ ) - 2. 뺐 뾿 26 Nu =vy (Wx= vy write kat at vy :- Xx+1 = + 크리트양 A Cs) +1 = 09 (+) Puri@ 2 ( -)1 =116979) vy 520 vy=520 = 3V 73 write kvL in outer loop (ho) Vab- 2x1 - Vy zo 1. Nab=a+vy =213 Vab = 5 V :: RN = Vab -1-52 (